OpenStudy (anonymous):

The Frostburg-Truth bus travels on a straight road from Frostburg Mall to Sojourner Truth Park. The mall is 2 miles east and 5 miles north of the City Center. The park is 5 miles west and 5 miles south of the Center. How far is it from the mall to the park to the nearest tenth of a mile?

OpenStudy (anonymous):

7.1 miles 13.5 miles 5.4 miles 12.2 miles

OpenStudy (anonymous):

a² = b² + c² a² = 7² + 10² a² = 49 + 100 a = √149 a = 12.20655

OpenStudy (anonymous):

circle the last one